OpenStudy (anonymous):

Using digits 0-9, find out how many 4 digit passcodes can be configured based on: A) the passcode can't start with a 0, a 1 or a 2 and no digits can be repeated. (I got 3,024-is this right?) B) The passcode must begin with an odd digit and end in an even digit (repetition OK)- I got 2,025 is this right?

OpenStudy (anonymous):

For A), I'd get \((10-3) \times 9P3\)

OpenStudy (anonymous):

For B) I get \(5\times 10P_r2\times 5\)

OpenStudy (anonymous):

So for A you got: 3,528? I don't understand how to do B the way you did it.

OpenStudy (anonymous):

\[ nP_rk = n^k \]

OpenStudy (anonymous):

It is just permutations with repetition.

OpenStudy (anonymous):

So B is 2250?

OpenStudy (anonymous):

For B I'm getting 2500

OpenStudy (anonymous):

The reason I multiply by 5 and start and end is because there are only 5 even/odd digits.

OpenStudy (anonymous):

I did it this way: 5*10*9*5=2250. (5*10!/(10-2)!*5)

OpenStudy (anonymous):

Did you get 3,528 for A?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

In B, can digits 2 and 3 be the same?

OpenStudy (anonymous):

Ok thank you- so I see you did factoring for B 5*10*10*5=2500

OpenStudy (anonymous):

Yes repetition is ok which is what I see you did with the 10*10 part.